# Are penalty shootouts better than a coin toss? Evidence from European football

### v2

@@ 5 Concluding remarks
Our paper has aimed to predict success in penalty shootouts in top European football during the last 25 years. We have considered several possible factors, such as the shooting order, the field of the match, and psychological momentum, but none of them affects the probability of winning. First in the literature, team strength is measured by Elo ratings instead of the division of the team or betting odds. This has led to a surprising finding: even though Elo ratings can forecast results in the UEFA Champions League, they entirely lose their predictive power if the qualification is decided by penalty shootouts.
These results could have important policy implications. First, if a drawn match is followed by penalty shootouts where the winner receives two points instead of one such as in the 2025 Leagues Cup, stronger teams would usually have more powerful incentives to attack since a draw means half extra points for both teams in expected value. Second, abolishing extra time, an intervention recently contemplated by UEFA, would favour weaker teams and improve competitive balance.
@@ Acknowledgements
We are grateful to Alex Krumer and Fabian Wunderlich for useful comments.
The research was supported by the National Research, Development and Innovation Office under Grants FK 145838 and PD 153835, and by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences.
@@ A.2 Supplementary tables
Table A.1: Results of binomial tests: first-mover advantage,
home advantage, momentum from the 2000/01 season to 2025
Potential success factor Wins Loses Total Win ratio $p$-value First-mover 143 125 268 53.4% 0.299 Home (extended data) 118 137 255 46.3% 0.260 Home (restricted data) 111 123 234 47.4% 0.472 Comeback (all goals) 72 75 147 49.0% 0.869 Comeback (last goal) 121 110 231 52.4% 0.511 Two-sided binomial tests. Significance: * $p<5\%$; ** $p<1\%$; *** $p<0.1\%$.
Table A.2: Results of binomial tests: team strength from the 2000/01 season to 2025
Scenario Favourite wins Favourite loses Total Win ratio $p$-value Elo difference >0>0 125 141 266 47.0% 0.358 Elo difference >25>25 115 127 242 47.5% 0.480 Elo difference >50>50 101 112 213 47.4% 0.493 Elo difference >75>75 84 95 179 46.9% 0.455 Elo difference >100>100 73 81 154 47.4% 0.573 Two-sided binomial tests. Significance: * $p<5\%$; ** $p<1\%$; *** $p<0.1\%$.
Table A.3: Results of binomial tests: first-mover advantage, home advantage,
momentum from the 2000/01 season to the 2019/20 season
Potential success factor Wins Loses Total Win ratio $p$-value First-mover 72 57 129 55.8% 0.218 Home (extended data) 64 58 122 52.5% 0.651 Home (restricted data) 64 57 121 52.9% 0.586 Comeback (all goals) 40 35 75 53.3% 0.644 Comeback (last goal) 53 55 108 49.1% 0.923 Two-sided binomial tests. Significance: * $p<5\%$; ** $p<1\%$; *** $p<0.1\%$.
@@
from the 2000/01 season to the 2019/20 season
Scenario Favourite wins Favourite loses Total Win ratio $p$-value Elo difference >0>0 60 69 129 46.5% 0.481 Elo difference >25>25 56 64 120 46.7% 0.523 Elo difference >50>50 50 58 108 46.3% 0.501 Elo difference >75>75 42 50 92 45.7% 0.466 Elo difference >100>100 36 40 76 47.4% 0.731 Two-sided binomial tests. Significance: * $p<5\%$; ** $p<1\%$; *** $p<0.1\%$.
Table A.5: Results of binomial tests: first-mover advantage,
home advantage, momentum from the 2020/21 season to 2025
Potential success factor Wins Loses Total Win ratio $p$-value First-mover 71 68 139 51.1% 0.865 Home (extended data) 54 79 133 40.6% 0.037* Home (restricted data) 47 66 113 41.6% 0.090 Comeback (all goals) 32 40 72 44.4% 0.410 Comeback (last goal) 68 55 123 55.3% 0.279 Two-sided binomial tests. Significance: * $p<5\%$; ** $p<1\%$; *** $p<0.1\%$.
Table A.6: Results of binomial tests: team strength from the 2020/21 season to 2025
Scenario Favourite wins Favourite loses Total Win ratio $p$-value Elo difference >0>0 65 72 137 47.4% 0.608 Elo difference >25>25 59 63 122 48.4% 0.786 Elo difference >50>50 51 54 105 48.6% 0.845 Elo difference >75>75 42 45 87 48.3% 0.830 Elo difference >100>100 37 41 78 47.4% 0.734 Two-sided binomial tests. Significance: * $p<5\%$; ** $p<1\%$; *** $p<0.1\%$.

### v3

@@ 6 Concluding remarks
Our paper has aimed to predict success in penalty shootouts in top European football during the last 25 years. We have considered several possible factors, such as the shooting order, the field of the match, and psychological momentum, but none of them affects the probability of winning significantly. First in the literature, team strength is measured by Elo ratings instead of the division of the team or betting odds. This has led to a somewhat surprising finding: even though Elo ratings can forecast results in the UEFA Champions League relatively well, they seem to lose their predictive power if the qualification is decided by penalty shootouts.
These results could have important policy implications. First, a drawn match can advance to a penalty shootout where the winner receives two points instead of one in order to increase attendance and prompt players to train more for penalty shootouts, which may be advantageous for the national team (Palacios-Huerta,, 2014). This rule was used in the 1988/89 season of the Argentinian League, and is currently applied in the Leagues Cup, an international club competition in North America, since 2023. If the outcome of a penalty shootout does not differ from a coin toss, stronger teams would have more powerful incentives to attack since a draw means half extra points for both teams in expected value. Second, abolishing extra time, an intervention recently contemplated by UEFA, would favour weaker teams and improve competitive balance.
@@ Acknowledgements
We are grateful to Alex Krumer and Fabian Wunderlich for their useful comments.
Two anonymous reviewers provided valuable remarks on an earlier draft.
The research was supported by the National Research, Development and Innovation Office under Grants Advanced 152220, FK 145838 and PD 153835, and by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences.
@@ A.2 Supplementary tables
Table A.1: Results of binomial tests: first-mover advantage, home advantage, momentum
Period 2000/01–2025 2000/01–2019/20 2020/21–2025 Potential success factor Total Ratio $p$-value Total Ratio $p$-value Total Ratio $p$-value First-mover 268 53.4% 0.299 129 55.8% 0.218 139 51.1% 0.865 Home (extended data) 255 46.3% 0.260 122 52.5% 0.651 133 40.6% 0.037* Home (restricted data) 234 47.4% 0.472 121 52.9% 0.586 113 41.6% 0.090 Comeback (all goals) 147 49.0% 0.869 75 53.3% 0.644 72 44.4% 0.410 Comeback (last goal) 231 54.5% 0.188 108 51.9% 0.773 123 56.9% 0.149 Two-sided binomial tests. Significance: * $p<5\%$; ** $p<1\%$; *** $p<0.1\%$. The column Total (Ratio) shows the number (proportion) of penalty shootouts won by the team indicated in the row.
@@
Period 2000/01–2025 2000/01–2019/20 2020/21–2025 Scenario Total Ratio $p$-value Total Ratio $p$-value Total Ratio $p$-value Elo difference >0>0 266 47.0% 0.358 129 46.5% 0.481 137 47.4% 0.608 Elo difference >25>25 242 47.5% 0.480 120 46.7% 0.523 122 48.4% 0.786 Elo difference >50>50 213 47.4% 0.493 108 46.3% 0.501 105 48.6% 0.845 Elo difference >75>75 179 46.9% 0.455 92 45.7% 0.466 87 48.3% 0.830 Elo difference >100>100 154 47.4% 0.573 76 47.4% 0.731 78 47.4% 0.734 Two-sided binomial tests. Significance: * $p<5\%$; ** $p<1\%$; *** $p<0.1\%$. The column Total (Ratio) shows the number (proportion) of penalty shootouts won by the favourite team.
Table A.3: Results of binomial tests in different tournament stages:
first-mover advantage, home advantage, momentum
Tournament phase Qualification Knockout Potential success factor Total Ratio $p$-value Total Ratio $p$-value First-mover 207 52.2% 0.578 61 57.4% 0.306 Home (extended data) 204 46.1% 0.294 51 47.1% 0.780 Home (restricted data) 184 47.3% 0.507 50 48.0% 0.888 Comeback (all goals) 108 49.1% 0.923 39 48.7% 1.000 Comeback (last goal) 176 53.4% 0.407 55 58.2% 0.281 Two-sided binomial tests. Significance: * $p<5\%$; ** $p<1\%$; *** $p<0.1\%$. The column Total (Ratio) shows the number (proportion) of penalty shootouts won by the team indicated in the row.
Table A.4: Results of binomial tests in different tournament stages: team strength
Tournament phase Qualification Knockout Scenario Total Ratio $p$-value Total Ratio $p$-value Elo difference >0>0 205 48.3% 0.675 61 42.6% 0.306 Elo difference >25>25 188 49.5% 0.942 54 40.7% 0.220 Elo difference >50>50 165 49.1% 0.876 48 41.7% 0.312 Elo difference >75>75 145 46.9% 0.507 34 47.1% 0.864 Elo difference >100>100 134 47.0% 0.546 20 50.0% 1.000 Two-sided binomial tests. Significance: * $p<5\%$; ** $p<1\%$; *** $p<0.1\%$. The column Total (Ratio) shows the number (proportion) of penalty shootouts won by the favourite team.